How much electricity is requirred to produce 40g Al from molten Al2O3?

Grade:12

2 Answers

Mayank
4 years ago
we have W=MQ/xF where M = atomic mass of Al Q=charge in columb x=valency factor F=faraday’s constant=96500
 
  now we have 40=(27*Q)/(3*96500)       // note: valency factor of Al is 3
  putting values as above we get Q=428,888.888 C
 
Khimraj
4 years ago
electricity  requirred to produce 27g Al2O3 = 3F
electricity  requirred to produce 40g Al2O3 = 3*40/27F = 4.44 F
Hope it clears........................